# EvalNet: A Practical Toolchain for Generation and Analysis of Extreme-Scale Interconnects

### v3

@@ 7.1. Which Simulator to Use?
LogGOPSim (Hoefler et al., 2017) 1024 (Hoefler et al., 2017) SM
htsim (Raiciu et al., 2010) 128 (Raiciu et al., 2010) SM
NS2 (Issariyakul and Hossain, 2009) $<$1,000 SM
booksim (Jiang et al., 2013) 10,000 (Besta and Hoefler, 2014) SM
FOGSim (Garcıa et al., 2014) 16,512 (Benito et al., 2015) SM
OMNet++(Varga and Hornig, 2008) 100,000 (Gamer and Mayer, 2009) SM
NS3 (of Washington NS-3 Consortium, 2024) 5,000 (Pelkey and Riley, 2011) DM
NS4 (Fan et al., 2017) N.A. DM
BigSim (Zheng et al., 2004) 46,656 (Jain et al., 2016) DM
SST (Rodrigues et al., 2016) 110,592 (Groves et al., 2016) DM
xSim (Böhm and Engelmann, 2011) 2,097,152 (Engelmann, 2014) DM
ROSS/CODES (Cope et al., 2011) 1,000,000 (Wolfe et al., 2016) (Slim Fly), 50,000,000 (Mubarak et al., 2012) (Dragonfly) DM
@@
Table 2. The comparison of available packet-level network simulators. “Scalability” is the largest size of a topology (simulated with a specific simulator) that we were able to find in the literature. “Design” includes the details of the compute platform used for simulations in the “Scalability” column: “SM” (shared-memory design), “DM” (distributed-memory design).
@@
packet (per flow) 10-1000 $10^{3}$$10^{10}$ (3) (5) 110 106,480,000 (3) (5) 110 104,742,000 (3) (5) 110 123,596,000 (3)
Table 3. Estimation of the number of elements of a data center Slim Fly simulation. The numbers and thus required resources vary widely depending on a chosen configuration. “Range” lists parameter bounds considered in this work, in the following columns we show numbers for example configurations. Memory estimates are based on htsim. The chosen workload parameters keep the simulations feasible on a laptop (configuration details are in 7.3). 1 Negligible for relevant configurations. 2 Data depends on the routing scheme. 3 Objects are transient and accounted for in the flow counts. 4 Inter-router cables + server links. 5 The average over the flow size distribution, excluding re-transmissions.
7.2. Can We Simulate Large Networks on Low-Cost HW?
@@
Number of Simulation Elements We estimate the number of elements of a simulated topology (e.g., servers) and of a simulated workload (e.g., flows which require flow control state). The results of the analysis are in Table 3. The number of network elements does not pose the most serious scalability problems. Instead, the number and corresponding size of elements related to workloads dominates the memory usage and the running time.
Memory and Time Requirements The total memory usage and simulation time depend on the simulation software and on granularity, i.e., what elements of the simulated workload are explicitly stored and simulated. We estimate these numbers in Table 3, which also shows what may not be feasible in the largest-scale simulations. Specifically, we cannot simulate large workloads on large networks, and must not store any per-packet state.
The lowest required amount of memory is determined by the number of simulated flows, because we want to simulate end-to-end flow control for each flow. Thus, the flow control algorithm state for each flow needs to be stored. We also account for the packets in flight on a per-flow basis, since the memory occupied by packets is tightly coupled to the number of concurrent flows. Memory consumed is about 2kB/flow plus 600B/path, which could probably be further reduced. To make large simulations feasible, it is crucial to keep the per-flow overhead low. Special attention needs to be paid to monitoring and debugging code, which can easily dominate the per-flow state. For very large networks, the quadratically increasing size of the routing tables starts to dominate, but as Table 3 shows, this was not a problem for our simulations.
@@
7.3. How to Simulate Really Large Networks on Low-Cost HW?
Finally, we show how to use htsim (Raiciu et al., 2010) and Booksim2 (Jiang et al., 2013) to enable more scalability. In Booksim2, the key limiting factor that we identified was processing related to single packets. This processing is usually related to gathering statistics, which is not relevant for the actual simulation outcomes. In Booksim2, eliminating this processing was the main enabler for being able to simulate networks with 10,000 endpoints and more. In htsim, on the contrary, the main bottleneck was the storage related to flows. Specifically, we had to add a more efficient, routing table based scheme for arbitrary topologies. We do not use the provided logging solution, which tends to produce too much output at the simulation scales that we consider. Instead, the statistics of each flow are printed in text format to standard output when the flow finishes. Another scalability obstacle was the net_paths structure present in the sample programs, a preallocated $N^{2}$ size matrix of routes, which would only be sparsely populated for permutation workloads and would dominate memory use for large networks.
7.4. Representative Simulation Results
We simulate networks with $\approx$10k, $\approx$100k, and $\approx$1M servers using a laptop with 16GB RAM and an Intel Core i7-8550U CPU; each run takes 1–3 hours. Networks are 5$\times$ oversubscribed, as in configurations studied in (Kassing et al., 2017), and include topologies larger than those used in prior large-scale SF simulations. Due to the cost of simulating large networks, we capture only a few milliseconds of activity—sufficient for reliable relative performance comparisons. We focus on flow completion time (FCT) distributions as a function of flow size, using the pFabric web search distribution (average $\approx$1MB) and injection rates of 40–60 flows/server/s. Results are shown as mean, 10th, and 99th percentiles. To mitigate artifacts from the fixed injection window, we carefully interpret short and long flows.
Figure 7. Results of large-scale simulations.
@@ 7.4. Representative Simulation Results
[Comparison of the Jellyfish and Slim Fly topologies]The figure shows three different plots comparing various aspects of the Jellyfish and Slim Fly topologies. The first plot presents the flow completion time with similar parameters as in the previous figure: Jellyfish and Slim are identical for the 10% percentile and the mean, but Jellyfish has a smaller completion time for the 99% percentile. The second subfigure plots the distribution as a histogram for 10.000, 100.000 and 1.000.000 servers for long flows (200KB) and an arrival rate of 50/s. Jellyfish exhibits a better tail latency. The third plot compares the flow completion time with respect to the arrival rate (ranging from 40 to 60) of long flows and the number of servers (10.000, 100.000 and 1.000.000). Both topologies exhibit the same mean performance, but Jellyfish is faster for the 10% and 90% percentile.
As an example, we analyze 10k, 100k, and 1M server simulations, see Figure 7. The left plot shows FCT as a function of size. For long flows ($v>200$kB), FCT is limited by bandwidth, while for shorter flows, latency dominates. The middle plot illustrates the distribution of FCTs for middle-size flows: here, the better tail behavior of the randomized Jellyfish topology is clearly visible. In the right plot, we show the impact of the flow arrival rate $\lambda$. Even a small increase in load causes a large degradation in FCT in the considered networks, due to the used oversubscription. The 1M server simulation is less affected due to the limited warmup time.
@@
8. COMPREHENSIVE ANALYSIS & TAKEAWAYS
@@
[…]
EvalNet enables multifaceted analysis of numerous network topologies. We list the considered networks in Table 4, and provide an assessment of their important properties (performance, cost & power consumption, path diversity), obtained using EvalNet. The table compiles the results described in the previous sections of this paper, as well as further results illustrated in the Appendix in Figures 856.
8.1. General Observations & Patterns
@@
8.2. Example Detailed Insights
@@
9. CONCLUSION
In this work, we introduce EvalNet, a comprehensive toolchain for the generation and in-depth analysis of extreme-scale interconnection networks, with a strong focus on path diversity as a core enabler of performance and scalability. EvalNet integrates topology generation, fine-grained path diversity analysis, performance modeling, and cost evaluation into a unified and extensible framework. Using EvalNet, we conduct the widest comparative study to date across 25 network topologies, revealing how shortest- and non-shortest-path diversity relate to key performance and deployment metrics such as latency, throughput, cost, and power. Our analysis shows that the ability to exploit non-minimal multipathing is essential for maintaining high throughput in low-diameter, scalable networks. EvalNet interfaces seamlessly with established simulators and serves as a practical and versatile tool for guiding future innovations in large-scale network design in data centers and supercomputers.
Acknowledgments
We thank Haoran Zhao for his code contributions. We thank Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). This project was supported by JSPS KAKENHI Grant Number JP19H04119.
@@ Appendix A Appendix
Figure 8. Shortest path plot for Kautz and Arrangement graphs. Kautz graphs have lower shortest paths compared to their Jellyfish equivalent for small networks. They also have lower shortest paths when compared to a small Arrangement network. A small Kautz graph has only a diameter of 3, while the Arrangement graph has already a diameter $\geq$ 5. The analysis of the Arrangement graphs compared to their Jellyfish equilibrium, which is a Jellyfish topology built with the same hardware, looks fairly similar. The Jellyfish equilibrium has less shortest paths of higher lengths $l$ = 5. The amount of shortest paths of low length (between 1 and 3) gets smaller, the bigger the network gets for all topologies. For Kautz graphs, these drop-offs are much more pronounced. There are about 100 times fewer shortest paths of lengths 1 when we compare a small Kautz graph with a large Kautz graph. In comparison, Arrangement networks have a drop-off between 5 and 10 times.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 9. Multiplicity plot for Kautz and Arrangement graphs. For Arrangement networks, the shortest path density first decreases, then increases. In contrast, Kautz graphs have about the same amount of path multiplicity, no matter the value of $n_{min}$. In the large Kautz network, we can see a significant drop off in the amount of multiplicity. This drop off is less significant for $K_{6,n}$. For Arrangement networks, these numbers of multiplicities stay about the same. There is only a small shift, for $n_{min}$ = 1 the number of shortest path multiplicities sinks for higher N, but for $n_{min}$ = 4 or 5 they only sink slightly. Kautz graphs have a gap in the multiplicity counts. They either have a low amount (1 or 2) of shortest paths between any two nodes or a lot of overlap ($\geq$ 4) in the multiplicity. The analysis of the Jellyfish equivalent looks similar to the Arrangement networks. The Jellyfish equivalent of the Kautz graphs on the other hand do not have this gap in the multiplicity. For $K_{4,n}$ and $K_{5,n}$ the drop off in the multiplicity count can also be seen in the Jellyfish equivalent for N $\approx$ 100,000. This cannot be seen for the Jellyfish equilibrium for $K_{6,n}$.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 10. Shortest path plot for mesh topology of different dimensions. The number of shortest paths increases while $l_{min}$ increases. For bigger networks, there are fewer shortest paths with a small value. The Jellyfish equivalent analysis looks fairly similar compared to the meshes. Only for $l_{min}$ = 5, the Jellyfish network with higher N does not strictly have fewer amounts of shortest paths.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 11. Multiplicity plot for meshes of different dimensions. As expected, the analysis looks fairly similar across all dimensions. For all sizes and all dimensions, the multiplicity count is the highest for $n_{min}\geq$ 5. For bigger network sizes, the amount of shortest path multiplicity decreases with roughly the same factor, independent of $n_{min}$. In contrast, the Jellyfish equivalent networks have almost no difference in the multiplicity count for $n_{min}$ = 1. For bigger values of $n_{min}$, the shortest path multiplicity drops even more compared to the original mesh, when we increase the network size.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 12. Shortest path plot for the 4D Express Meshes with different gap sizes. We choose the lowest g, such that for each Express Mesh 4DeMESHx we have at least x additional express connections per dimension. As one can expect, the Express Meshes have lower shortest paths than the original mesh due to their express connections. For some network sizes, the Express Meshes have only a diameter of 4, while the standard 4D mesh has a diameter of $\geq$ 5. While the number of shortest paths increase for bigger $l$ in the Express Meshes, their equivalent Jellyfishes don’t show a similar pattern. They increase first before they decrease for the biggest value of $l$ where there still exists a shortest path.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 13. Multiplicity plot for Express Meshes of dimension 4. For small and medium-sized networks, there is a gap in the multiplicity counts. The length of the shortest path for any two nodes exists either once or twice, or $\geq$ 5 times. For large Express Meshes, this gap is not visible anymore. Express Meshes keep the high multiplicity of shortest paths, even for higher N, due to their express connections, especially for lower gap sizes. For all network sizes, the Express Meshes have the highest shortest paths multiplicity count for $n_{min}\geq$ 5. There is a big difference in the multiplicity of the meshes compared to the equivalent Jellyfish. In contrast to the base Express Meshes, the Jellyfish equilibriums are well distributed along all values for $n_{min}$ for small and medium networks.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 14. Shortest path plot for 6D Express Meshes with different gap sizes. We chose the lowest g, such that for each Express Mesh 6DeMESHx we have at least x additional express connections per dimension. The shortest path analysis of these Express Meshes look similar to the analysis of their 4D counterpart. Express connections do not allow the 6D graph to become a diameter $\leq$ 5 network, since any path only traverses over one dimension. The Jellyfish equilibrium is also fairly similar when comparing the 6D to the 4D analysis. The amount of shortest paths grows as $l$ grows, before dropping down for some outliers that have a shortest path length of 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 15. Multiplicity plot for Express Meshes of dimension 6. We can see again the gap in the amount of shortest path multiplicity. There either exist very few paths of length equal to the shortest path between two nodes, or there exists a lot of them (more than 4). Only the large Express Mesh with two additional express connections per dimension does not show this gap. Same as in the 4D case, most shortest paths have a multiplicity $\geq$ 5. In case of the Jellyfish equilibrium, the shortest path multiplicity count drops at higher $n_{min}$. Just for $n_{min}$ = 5, the multiplicity count increases, as they have as well the most shortest paths with multiplicity $\geq$ 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 16. Shortest path plot for the Tofu topology, 6D torus and 6D mesh. As we have already seen in the analysis of the meshes, the amount of shortest paths sinks as the network size gets bigger. The bigger $l$ gets, the more shortest paths exist of length $l$. All networks show similar patterns. Tofu has more shortest paths $l_{min}\leq$ 5 compared to all the other meshes for small networks, but is worse for bigger networks. The same is true when we compare Tofu to the equivalent Jellyfish.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 17. Multiplicity plot for the Tofu network, 6D torus and 6D mesh. Similar as for the meshes, all the topologies have the most multiplicity count where $n_{min}\geq$ 5. The Tofu network has even the highest multiplicity count for $n_{min}\geq$ 5 for all small networks. In medium and large networks, Tofu does not have the most multiplicity count anymore. The bigger the network sizes are, the smaller the multiplicity counts get. In the analysis of the torus, we can find the already mentioned gap in the shortest path multiplicities for small and medium-size networks. Tofu has this gap only for small networks. The Jellyfish equilibrium of the Tofu network is well distributed along all values for $n_{min}$ for small and medium networks. For the large network, there is a significant drop-off and the multiplicity is about 30 times smaller for $n_{min}\geq$ 5 compared to the small or medium network.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 18. Shortest path plot for PolarStar compared to Dragonfly and Slim Fly. We can clearly see that all are diameter $\leq$ 3 topologies. All of the topologies have in general lower smallest paths compared to their equivalent Jellyfish graph. Slim Fly has only shortest paths of length 2. All the others look fairly similar, having fewer shortest paths with $l_{min}$ = 1 compared to $l_{min}$ = 2 or 3. The amount of shortest paths of length 1 and 2 decreases as the network size increases. The shortest paths of length 3 slightly increase though. There is almost no difference for the analysis between PolarStar with BDF or Paley subgraphs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 19. Multiplicity plot for PolarStar, Dragonfly and Slim Fly. Slim Fly has the least multiple shortest paths of all topologies. Only for medium and large networks, there are pairs of nodes with shortest path multiplicity bigger than 3. All topologies have the most shortest paths only once between any two nodes. PolarStar with BDF subgraphs has only a few shortest path multiplicity count for $n_{min}$ = 3 or 4, while PolarStar with the Paley subgraphs has a well-distributed count for all values of $n_{min}$.
[…]
[Main findings are in the caption.]Main findings are in the caption.

### v4

@@ VII-A Which Simulator to Use?
TABLE II: The comparison of available packet-level network simulators. “Scalability” is the largest size of a topology (simulated with a specific simulator) that we were able to find in the literature. “Design” includes the details of the compute platform used for simulations in the “Scalability” column: “SM” (shared-memory design), “DM” (distributed-memory design).
@@
LogGOPSim [46] 1024 [46] SM
htsim [91] 128 [91] SM
NS2 [53] <<1,000 SM
booksim [56] 10,000 [15] SM
FOGSim [39] 16,512 [12] SM
OMNet++[103] 100,000 [38] SM
NS3 [99] 5,000 [82] DM
NS4 [35] N.A. DM
BigSim [110] 46,656 [55] DM
SST [92] 110,592 [41] DM
xSim [19] 2,097,152 [33] DM
ROSS/CODES [27] 1,000,000 [105] (Slim Fly), 50,000,000 [77] (Dragonfly) DM
@@
TABLE III: Estimation of the number of elements of a data center Slim Fly simulation. The numbers and thus required resources vary widely depending on a chosen configuration. “Range” lists parameter bounds considered in this work, in the following columns we show numbers for example configurations. Memory estimates are based on htsim. The chosen workload parameters keep the simulations feasible on a laptop (configuration details are in Section VII-C). 1Negligible for relevant configurations. 2Data depends on the routing scheme. 3Objects are transient and accounted for in the flow counts. 4Inter-router cables + server links. 5The average over the flow size distribution, excluding re-transmissions.
@@
packet (per flow) 10-1000 $^{3}1010$$^{10}$ (3) (5) 110 106,480,000 (3) (5) 110 104,742,000 (3) (5) 110 123,596,000 (3)
VII-B Can We Simulate Massive Networks on Low-Cost HW?
@@
Number of Simulation Elements We estimate the number of elements of a simulated topology (e.g., servers) and of a simulated workload (e.g., flows which require flow control state). The results of the analysis are in Table III. The number of network elements does not pose the most serious scalability problems. Instead, the number and corresponding size of elements related to workloads dominates the memory usage and the running time.
Memory and Time Requirements The total memory usage and simulation time depend on the simulation software and on granularity, i.e., what elements of the simulated workload are explicitly stored and simulated. We estimate these numbers in Table III, which also shows what may not be feasible in the largest-scale simulations. Specifically, we cannot simulate large workloads on large networks, and must not store any per-packet state.
The lowest required amount of memory is determined by the number of simulated flows, because we want to simulate end-to-end flow control for each flow. Thus, the flow control algorithm state for each flow needs to be stored. We also account for the packets in flight on a per-flow basis, since the memory occupied by packets is tightly coupled to the number of concurrent flows. Memory consumed is about 2kB/flow plus 600B/path, which could probably be further reduced. To make large simulations feasible, it is crucial to keep the per-flow overhead low. Special attention needs to be paid to monitoring and debugging code, which can easily dominate the per-flow state. For very large networks, the quadratically increasing size of the routing tables starts to dominate, but as Table III shows, this was not a problem for our simulations.
@@
VII-C How to Simulate Massive Networks on Low-Cost HW?
Finally, we show how to use htsim and Booksim2 to enable more scalability. In Booksim2, the key limiting factor that we identified was processing related to single packets. This processing is usually related to gathering statistics, which is not relevant for the actual simulation outcomes. In Booksim2, eliminating this processing was the main enabler for being able to simulate networks with 10k to 1M endpoints and more. In htsim, on the contrary, the main bottleneck was the storage related to flows. Specifically, we had to add a more efficient, routing table based scheme for arbitrary topologies. We do not use the provided logging solution, which tends to produce too much output at the simulation scales that we consider. Instead, the statistics of each flow are printed in text format to standard output when the flow finishes. Another scalability obstacle was the net_paths structure present in the sample programs, a preallocated $N^{2}$ size matrix of routes, which would only be sparsely populated for permutation workloads and would dominate memory use for large networks.
VII-D Representative Simulation Results
@@
Figure 7: Results of large-scale simulations.
As an example, we analyze 10k, 100k, and 1M server simulations, see Figure 7. The left plot shows FCT as a function of size. For long flows (v>200v>200kB), FCT is limited by bandwidth, while for shorter flows, latency dominates. The middle plot illustrates the distribution of FCTs for middle-size flows: here, the better tail behavior of the randomized Jellyfish topology is clearly visible. In the right plot, we show the impact of the flow arrival rate $\lambda$. Even a small increase in load causes a large degradation in FCT in the considered networks, due to the used oversubscription. The 1M server simulation is less affected due to the limited warmup time.
@@
VIII COMPREHENSIVE ANALYSIS & TAKEAWAYS
TABLE IV: A general comparison of different network topologies. Lat.: latency. GlB.: global bandwidth (i.e., the saturation point for the random uniform traffic pattern). Ext.: extensibility (i.e., how far can we extend a given concrete network with new servers?). Note that fixed diameter networks have lowest extensibility, because they have strict upper bounds on how many new servers can be attached; the lower the diameter is, the lower this bound is. Contrarily, networks such as torus can attach arbitrarily many new servers, because their diameter grows to infinity. C&P: construction cost and static power consumption. They are assessed as proportional to the total number of ports in a network, for a fixed network size NN (i.e., for fixed $N$, networks with higher radix tend to have higher cost and power consumption). MS: diversity of shortest paths. It is assessed as a weighted average of shortest path diversities at different lengths. MnS: diversity of non-shortest paths. It is assessed as a weighted average of non-shortest path diversities at different lengths. The battery symbols serve as a rudimentary measure of relative comparison between networks. [0]: worst, [1]: bad, [2]: medium, [3]: good, [4]: excellent. They always have a positive meaning, e.g., “ [4]” used for cost means that the cost of a given network is very low compared to other networks.
@@ VIII COMPREHENSIVE ANALYSIS & TAKEAWAYS
Slim Fly [15] [4] [4] [0] [3] [1] [4]
PolarFly [65] [4] [4] [0] [4] [0] [4]
Xpander [100] [3] [4] [1] [2] [1] [3]
BundleFly [68] [3] [3] [1] [2] [1] [3]
Dragonfly [63] [3] [3] [1] [2] [0] [2]
Cascade Dragonfly [34] [2] [3] [1] [2] [1] [1]
Jellyfish [94] [2] [2] [1] [2] [1] [3]
Mesh [0] [0] [3] [0] [0] [1]
Torus 2D [0] [0] [3] [0] [0] [1]
Torus 3D [1] [1] [3] [1] [1] [2]
Torus 4D [2] [2] [3] [1] [2] [3]
Torus 5D [3] [3] [4] [2] [3] [3]
Torus 6D [3] [3] [4] [2] [3] [4]
Hypercube [3] [3] [4] [2] [3] [4]
HyperX (2-dimensional) [1] [4] [4] [1] [2] [3] [4]
HyperX (3-dimensional) [1] [3] [4] [2] [2] [3] [4]
Express Mesh [54] [2] [3] [3] [3] [2] [3]
Flattened Butterfly [64] [4] [4] [1] [2] [3] [4]
Fat Tree (2-level) [69] [4] [4] [1] [3] [4] [0]
Fat Tree (3-level) [69] [3] [2] [2] [2] [4] [0]
k-ary n-tree [84] [3] [4] [4] [2] [4] [0]
eXtended Generalized Fat Tree [79] [3] [3] [4] [2] [4] [0]
Orthogonal Fat-Trees [60] [4] [4] [1] [4] [4] [0]
Multi-Layer Full-Mesh [60] [4] [3] [1] [2] [3] [2]
Undirected Kautz [70] [2] [2] [1] [2] [1] [2]
Arrangement Network [30] [1] [2] [1] [2] [0] [1]
EvalNet enables multifaceted analysis of numerous network topologies. We list the considered networks in Table IV, and provide an assessment of their important properties (performance, cost & power consumption, path diversity), obtained using EvalNet. The table compiles the results described in the previous sections of this paper, as well as further results illustrated in the Appendix in Figures 856.
VIII-A General Observations & Patterns
@@
VIII-B Example Detailed Insights
@@
IX CONCLUSION
We introduce EvalNet, the first toolchain that enables comprehensive generation and in-depth analysis of extreme-scale interconnection networks, with a strong focus on path diversity as a core enabler of performance and scalability. EvalNet integrates topology generation, fine-grained path diversity analysis, performance modeling, and cost evaluation into a unified and extensible framework. Using EvalNet, we conduct the widest comparative study to date across 25 network topologies, revealing how shortest- and non-shortest-path diversity relate to key performance and deployment metrics such as latency, throughput, cost, and power. Our analysis shows that the ability to exploit non-minimal multipathing is essential for maintaining high throughput in low-diameter, scalable networks. EvalNet interfaces seamlessly with established simulators and serves as a practical and versatile tool for guiding future innovations in large-scale network design in data centers and supercomputers.
Acknowledgment
We thank Ankit Singla for help during the early stages and Haoran Zhao for his code contributions. We thank Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines. We thank Timo Schneider for help with infrastructure at SPCL. This paper has been assigned the LANL identification number LA-UR-26-20454. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). This project was supported by JSPS KAKENHI Grant Number JP19H04119. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grants no. PLG/2024/017103 and PLG/2025/018259.
@@ Appendix A Appendix
Figure 8: Shortest path plot for Kautz and Arrangement graphs. Kautz graphs have lower shortest paths compared to their Jellyfish equivalent for small networks. They also have lower shortest paths when compared to a small Arrangement network. A small Kautz graph has only a diameter of 3, while the Arrangement graph has already a diameter $\geq$ 5. The analysis of the Arrangement graphs compared to their Jellyfish equilibrium, which is a Jellyfish topology built with the same hardware, looks fairly similar. The Jellyfish equilibrium has less shortest paths of higher lengths $l$ = 5. The amount of shortest paths of low length (between 1 and 3) gets smaller, the bigger the network gets for all topologies. For Kautz graphs, these drop-offs are much more pronounced. There are about 100 times fewer shortest paths of lengths 1 when we compare a small Kautz graph with a large Kautz graph. In comparison, Arrangement networks have a drop-off between 5 and 10 times.
@@
Figure 9: Multiplicity plot for Kautz and Arrangement graphs. For Arrangement networks, the shortest path density first decreases, then increases. In contrast, Kautz graphs have about the same amount of path multiplicity, no matter the value of $n_{min}$. In the large Kautz network, we can see a significant drop off in the amount of multiplicity. This drop off is less significant for $K_{6,n}$. For Arrangement networks, these numbers of multiplicities stay about the same. There is only a small shift, for $n_{min}$ = 1 the number of shortest path multiplicities sinks for higher N, but for $n_{min}$ = 4 or 5 they only sink slightly. Kautz graphs have a gap in the multiplicity counts. They either have a low amount (1 or 2) of shortest paths between any two nodes or a lot of overlap ($\geq$ 4) in the multiplicity. The analysis of the Jellyfish equivalent looks similar to the Arrangement networks. The Jellyfish equivalent of the Kautz graphs on the other hand do not have this gap in the multiplicity. For $K_{4,n}$ and $K_{5,n}$ the drop off in the multiplicity count can also be seen in the Jellyfish equivalent for N $\approx$ 100,000. This cannot be seen for the Jellyfish equilibrium for $K_{6,n}$.
@@
Figure 10: Shortest path plot for mesh topology of different dimensions. The number of shortest paths increases while $l_{min}$ increases. For bigger networks, there are fewer shortest paths with a small value. The Jellyfish equivalent analysis looks fairly similar compared to the meshes. Only for $l_{min}$ = 5, the Jellyfish network with higher N does not strictly have fewer amounts of shortest paths.
@@
Figure 11: Multiplicity plot for meshes of different dimensions. As expected, the analysis looks fairly similar across all dimensions. For all sizes and all dimensions, the multiplicity count is the highest for $n_{min}\geq$ 5. For bigger network sizes, the amount of shortest path multiplicity decreases with roughly the same factor, independent of $n_{min}$. In contrast, the Jellyfish equivalent networks have almost no difference in the multiplicity count for $n_{min}$ = 1. For bigger values of $n_{min}$, the shortest path multiplicity drops even more compared to the original mesh, when we increase the network size.
@@
Figure 12: Shortest path plot for the 4D Express Meshes with different gap sizes. We choose the lowest g, such that for each Express Mesh 4DeMESHx we have at least x additional express connections per dimension. As one can expect, the Express Meshes have lower shortest paths than the original mesh due to their express connections. For some network sizes, the Express Meshes have only a diameter of 4, while the standard 4D mesh has a diameter of $\geq$ 5. While the number of shortest paths increase for bigger $l$ in the Express Meshes, their equivalent Jellyfishes don’t show a similar pattern. They increase first before they decrease for the biggest value of $l$ where there still exists a shortest path.
@@
Figure 13: Multiplicity plot for Express Meshes of dimension 4. For small and medium-sized networks, there is a gap in the multiplicity counts. The length of the shortest path for any two nodes exists either once or twice, or $\geq$ 5 times. For large Express Meshes, this gap is not visible anymore. Express Meshes keep the high multiplicity of shortest paths, even for higher N, due to their express connections, especially for lower gap sizes. For all network sizes, the Express Meshes have the highest shortest paths multiplicity count for $n_{min}\geq$ 5. There is a big difference in the multiplicity of the meshes compared to the equivalent Jellyfish. In contrast to the base Express Meshes, the Jellyfish equilibriums are well distributed along all values for $n_{min}$ for small and medium networks.
@@
Figure 14: Shortest path plot for 6D Express Meshes with different gap sizes. We chose the lowest g, such that for each Express Mesh 6DeMESHx we have at least x additional express connections per dimension. The shortest path analysis of these Express Meshes look similar to the analysis of their 4D counterpart. Express connections do not allow the 6D graph to become a diameter $\leq$ 5 network, since any path only traverses over one dimension. The Jellyfish equilibrium is also fairly similar when comparing the 6D to the 4D analysis. The amount of shortest paths grows as $l$ grows, before dropping down for some outliers that have a shortest path length of 5.
@@
Figure 15: Multiplicity plot for Express Meshes of dimension 6. We can see again the gap in the amount of shortest path multiplicity. There either exist very few paths of length equal to the shortest path between two nodes, or there exists a lot of them (more than 4). Only the large Express Mesh with two additional express connections per dimension does not show this gap. Same as in the 4D case, most shortest paths have a multiplicity $\geq$ 5. In case of the Jellyfish equilibrium, the shortest path multiplicity count drops at higher $n_{min}$. Just for $n_{min}$ = 5, the multiplicity count increases, as they have as well the most shortest paths with multiplicity $\geq$ 5.
@@
Figure 16: Shortest path plot for the Tofu topology, 6D torus and 6D mesh. As we have already seen in the analysis of the meshes, the amount of shortest paths sinks as the network size gets bigger. The bigger $l$ gets, the more shortest paths exist of length $l$. All networks show similar patterns. Tofu has more shortest paths $l_{min}\leq$ 5 compared to all the other meshes for small networks, but is worse for bigger networks. The same is true when we compare Tofu to the equivalent Jellyfish.
@@
Figure 17: Multiplicity plot for the Tofu network, 6D torus and 6D mesh. Similar as for the meshes, all the topologies have the most multiplicity count where $n_{min}\geq$ 5. The Tofu network has even the highest multiplicity count for $n_{min}\geq$ 5 for all small networks. In medium and large networks, Tofu does not have the most multiplicity count anymore. The bigger the network sizes are, the smaller the multiplicity counts get. In the analysis of the torus, we can find the already mentioned gap in the shortest path multiplicities for small and medium-size networks. Tofu has this gap only for small networks. The Jellyfish equilibrium of the Tofu network is well distributed along all values for $n_{min}$ for small and medium networks. For the large network, there is a significant drop-off and the multiplicity is about 30 times smaller for $n_{min}\geq$ 5 compared to the small or medium network.
@@
Figure 18: Shortest path plot for PolarStar compared to Dragonfly and Slim Fly. We can clearly see that all are diameter $\leq$ 3 topologies. All of the topologies have in general lower smallest paths compared to their equivalent Jellyfish graph. Slim Fly has only shortest paths of length 2. All the others look fairly similar, having fewer shortest paths with $l_{min}$ = 1 compared to $l_{min}$ = 2 or 3. The amount of shortest paths of length 1 and 2 decreases as the network size increases. The shortest paths of length 3 slightly increase though. There is almost no difference for the analysis between PolarStar with BDF or Paley subgraphs.
@@
Figure 19: Multiplicity plot for PolarStar, Dragonfly and Slim Fly. Slim Fly has the least multiple shortest paths of all topologies. Only for medium and large networks, there are pairs of nodes with shortest path multiplicity bigger than 3. All topologies have the most shortest paths only once between any two nodes. PolarStar with BDF subgraphs has only a few shortest path multiplicity count for $n_{min}$ = 3 or 4, while PolarStar with the Paley subgraphs has a well-distributed count for all values of $n_{min}$.